OpenStudy (anonymous):

A line passes through the point (5,-8) and has an x intercept of -3. Find the equation of the line in slope intercept form.

OpenStudy (anonymous):

y=mx+b form

OpenStudy (anonymous):

Apply the point slope form (y - y1) = m(x - x1) , where m is the slope and (x1, y1) is the point (5, -8)

OpenStudy (anonymous):

write y=mx+b with the known y intercept then plug in the point (instead of y and x ) and solve for slope.

OpenStudy (loser66):

x-intercept of -3 means y =0 at x =-3 or (-3,0) is another point. So far, you have 2 points (-3,0) and (5,-8). Who stops you write down the equation of the line?

OpenStudy (anonymous):

We know the equation has the pts: (5, -8) and (-3, 0) Now we get the amount of change in the y direction: 0 - (-8) = 8 And the amount of change in the x direction: -3 - 5 = -8 The slope of the line is a number that tells you how much y changes for each unit of x. Logically, this is: the amount y changes/the amount x changes slope = -8/8 = -1 so we have a line y = -1x + b, where b is some constant Plug in any point (I'm plugging in (-3,0): 0 = 3 + b b = -3 So we have: y = -1x - 3

OpenStudy (mathmale):

Hey, guys: Please help Alphabetical to solve this problem herself, with your guidance / hints / answers to her questions, etc. Thanks.

OpenStudy (anonymous):

Thanks for the help everyone. It is appreciated.